OpenStudy (anonymous):

2y-2x+2=0 y= how do i get the slope intercept?

OpenStudy (anonymous):

that is just a different form of the graphing equations and i really can't think of the name of it off the top of my head. basically make it look like y=mx+b......first +2y to both sides to get rid of the y on the left and you have 2x+2 = 2y then divide both sides by 2 to get y all by itself and you get x + 1 = y or y=x+1 so what is the slope if out of y=mx+b by looking at that equation

OpenStudy (anonymous):

ok i read that wrong you said slope intercept. do you mean the slope, the intercept, or slope intercept form. if you mean the form of the equations, that's y=mx+b and i wrote it for you. if you mean the intercept, that is b out of the y=mx+b equation.